i need help solving 21-17t+2t^2
Answer by stanbon(75887)

i need help solving 21-17t+2t^2
---
a = 2 ; b = -17 ; c = 21
---
Quadratic formula:
t = [17 +- sqrt(17^2-4*2*21)]/4
----
t = [17 +- sqrt(121)]/4
---
t = [17 +- 11]/4
---
t = (17+11)/4 or t = (17-11)/4
---
t = 7 or t = 3/2
----
Factors would be (7-t)(3-2t)
=============================
Cheers,
Stan H.